A rectangular room is 1.7 times as long as it is wide, and its perimeter is 28 meters. Find the dimension of the room.
The length is : ??? meters and the width is ??? meters.

Answer by ewatrrr(24785)


Hi
Let x and 1.7x represent the width and length respectively
P = 2*length + 2*width (Distance all away around)
Question states***
2(1.7x) + 2x = 28m
3.4x + 2x = 28m
5.4x = 28m
x = 28m/5.4
x = 5.2m, the width. The length = 8.8m (1.7*5.2)
Answers given 1n the nearest tenth of a meter
CHECKING our Answer
2*5.2 + 2*8.8 = 10.4 + 17.6 = 28m